FURTHER STUDIES OF POTASSIUM-BEARING MOLECULES : THE MILLIMETER-WAVE SPECTRUM OF KSH (X^{1}A′)

\underline{MATTHEW P. BUCCHINO}, Department of Chemistry and Biochemistry, University of Arizona, Tucson, Arizona 85721; JUSTIN P. YOUNG, PHIL M. SHERIDAN and DAVID EWING, Department of Chemistry, Canisius College, Buffalo, New York 14208; LUCY M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, Arizona 85721.

The pure rotational spectrum of KSH (X^{1}A′) has been measured using millimeter-wave direct-absorption techniques for the first time. KSH was synthesized by the reaction of potassium vapor and H_{2}S in the presence of a direct current discharge. Five rotational transitions in the frequency range of 260−300 GHz were recorded for both KSH and KSD in their ground electronic states. The K-ladder structure indicates C_{s} symmetry, and therefore a bent molecule. Accurate rotational and hyperfine constants have been determined from a global fit including previously measured microwave spectra. A comparison of KSH to other alkali metal hydrosulfides will be presented.